OpenStudy (anonymous):

Solve for x: 2 over 3 (x − 4) = 2x. 2 −2 −8 −4

OpenStudy (anonymous):

2/3 (x - 4) = 2x --- distribute through the parenthesis 2/3x - 8/3 = 2x --- multiply everything by common denominator of 3 to get rid of fractions 2x - 8 = 6x -- subtract 2x -8 = 4x -- divide by 4 -2 = x check.. 2/3 (-2 - 4) = 2 (-2) 2/3 (-6) = -4 -12/3 = -4 -4 = -4 (correct) so x = -2 questions ?

OpenStudy (anonymous):

How does 2/3(-6) became -12/3 ???